OpenStudy (anonymous):

Simplify log8 7 - log8 s + log8 t - log8 4

OpenStudy (anonymous):

Log8(7/8*8/4)=log8(7/4)

OpenStudy (unklerhaukus):

\[\log_b(x)+\log_b(y)=\log_b(xy)\] \[\log_b(x)-\log_b(y)=\log_b(x/y)\]

OpenStudy (anonymous):

7s/4t or would it be 7t/4s. I knew 8 would be the base, it's just the fraction that keeps messing me up

OpenStudy (unklerhaukus):

7/s * t/4 use a/b * c/d = ac/bd

OpenStudy (unklerhaukus):

7t/4s is righht

OpenStudy (anonymous):

alright, thank you :)

OpenStudy (unklerhaukus):

but dont forget the log_8 part

OpenStudy (anonymous):

okay, I already had that figured out before I asked